A 10cm×10cm×10cm wood block with a density of 650 kg/m3 floats in water.

Part A

What is the distance from the top of the block to the water if the water is fresh?

.


Part B

If it's seawater? Suppose that ρ=1030kg/m3.
Answer #1

(a).

density of fresh water is ρ_water = 1000 kg/m³

ΣF = 0

- (ρ_wood)g(V_wood) + (ρ_water)g(v_wood) = 0

(v_wood)/(V_wood) = (ρ_wood)/(ρ_water)

(10 * 10 * h)/(10 * 10 * 10) = (650)/(1000)

h = 6.5 cm

the distance from the top of the block to the water is 10 - 6.5 = 3.5 cm....Answer.




(b).

density of seawater is ρ_sea = 1030 kg/m³

(v_wood)/(V_wood) = (ρ_wood)/(ρ_sea)

(10 * 10 * h')/(10 * 10 * 10) = (650)/(1030)

h' = 6.31 cm

the distance from the top of the block to the water is 10 - 6.31 = 3.7 cm...Answer.

HOPE THIS HELPS!

Please thumbs up !